If the probability of an event is 0.3, the probability of it not happening is:
- 0
- 0.3
- 0.7
- 1.3
Show answer and solution
Correct answer: C. 0.7
P(not E) = 1 − 0.3 = 0.7.
Explore 75 free Statistics and Probability questions in UP Board Class 10 Mathematics. Open each question or exercise to read its content, reference answers and available explanation.
Practise Statistics and Probability with Quill →Showing questions 31–60 of 75.
Correct answer: C. 0.7
Correct answer: A. 0
Correct answer: D. 1/2
Continue with Statistics and Probability practice for UP Board Class 10. Read explanations as you go. No signup form needed to start.
Practise Statistics and Probability →